OpenStudy (anonymous):

Nathan had an infection, and his doctor wanted him to take penicillin. Because Nathan’s father and paternal grandfather were allergic to penicillin, Nathan had a 75% chance of having the same allergy. The doctor performed a skin test to see whether Nathan would react to it. The test is 98% accurate. What is the probability that Nathan is allergic to penicillin and the test predicts it?

OpenStudy (anonymous):

that would be 86.5% :))))

OpenStudy (here_to_help15):

Nathan has 0.75 chance of allergy 0.25 chance of no allergy test is 0.98 correct 0.02 incorrect Nathan is not allergic to penicillin and the test predicts it or in other words Nathan no allergy and test incorrect multiply the probabilities: Pr(no allergy) * Pr(incorrect) = 0.25*0.02 = 0.005

OpenStudy (anonymous):

I hate skin tests OMG

OpenStudy (here_to_help15):

Wait im wrong let me see

OpenStudy (anonymous):

0.005 0.015 0.245 0.735 These are the answer choices

OpenStudy (here_to_help15):

wait then i might be right : P

OpenStudy (anonymous):

I think you are @Here_to_Help15

OpenStudy (anonymous):

:))))))))))

OpenStudy (anonymous):

It was not .005

OpenStudy (anonymous):

Thanks. I've got it wrong like 5 times already. I tried :P

OpenStudy (anonymous):

0.245